OpenStudy (kaloemalia):

algebra help please !!! simplify -4(x-5)+3(a-7)

OpenStudy (kaloemalia):

@Michele_Laino @Nnesha

OpenStudy (cggurumanjunath):

use distributive property.

OpenStudy (kaloemalia):

okay, one second please!

OpenStudy (kaloemalia):

(x-20)-7(a-11) ???

OpenStudy (kaloemalia):

thats wrong huh?

OpenStudy (cggurumanjunath):

-4(x-5)=-4*x+(-4)*(-5)=-4x+20

OpenStudy (cggurumanjunath):

now can you use distributive property a(b+c) =a*b+a*c for 3(a-7)

OpenStudy (kaloemalia):

-4x+3a+1 ??

OpenStudy (kaloemalia):

-4x+3a-1

OpenStudy (cggurumanjunath):

:)

OpenStudy (kaloemalia):

the second one?

OpenStudy (cggurumanjunath):

perfect !

OpenStudy (kaloemalia):

thank you!

OpenStudy (cggurumanjunath):

you're always welcome. Tag me anytime .

OpenStudy (cggurumanjunath):

when you need any Math help !